# Semitoric Families on Pentagon Spaces

### v1

@@ 3.2.2. Rank one points
Recall that a point $p\in M$ has rank one if $dJ_{p},dH_{p}$ span a one-dimensional vector space. When $dJ_{p}\neq 0$, $p$ is not a fixed point of the $S^{1}$-action. If $S^{1}$ acts freely on $p$, we can apply symplectic reduction (at least locally) at the level $J(p)$ and study the type of singularity of $p]\in M^{\textrm{red},J(p)}=J^{-1}(J(p))/S^{1}$ of the function $H^{\textrm{red},J(p)}$ to which $H$ descends.
@@
[LFP23, Lemma 2.6] Let $(M,\omega,F=(J,H))$ be a 4-dimensional integrable system, where $J$ is the moment map of an effective Hamiltonian $S^{1}$-action on $M$. Let $p\in M$ be a point on which $S^{1}$ acts freely. Let $j=J(p)$. Then $p$ is a rank one singular point of $F$ if and only if $[p]$ is a singular point of $H^{\textrm{red},j}$. Moreover, $p$ is a non-degenerate singular point of $F$ if and only if $[p]$ is non-degenerate for $H^{\textrm{red},j}$ (in the Morse sense) and the type of $p$ is elliptic-regular (respectively hyperbolic-regular) if and only if $[p]$ is an elliptic (respectively hyperbolic) singular point of $H^{\textrm{red},j}$.
@@
The proof of the first claim is a similar computation to the proof of Lemma 3.5, so we leave it to the readers to verify. Since $X_{J}(0)=0$, the real span of $X_{J}(0),X_{H}(0)$ is the same as the real span of $X_{H}(0)$. Let $L$ denote this real span. Then $L^{\omega}={\mathbb{R}}(\frac{\partial}{\partial z_{1}},\frac{\partial}{\partial\overline{z}_{1}},X_{H}(0))$. Hence, we can use the equivalent classes of $\frac{\partial}{\partial z_{1}},\frac{\partial}{\partial\overline{z}_{1}}$ as a basis of $L^{\omega}/L$. Under this basis, we have the following matrix representation:
$\Omega=\begin{pmatrix}0&\frac{i}{2}\\ -\frac{i}{2}&0\end{pmatrix}\hskip 56.9055pt\operatorname{Hess}(J)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$
Hence, $A_{1,0}=\Omega^{-1}\operatorname{Hess}(J)$ has eigenvalues $\pm 2i$, so by definition, $(0,0)$ is an elliptic-regular point of $(J,H)$. ∎
@@
Let $(M,\omega,(J,H))$ be a 4-dimensional integrable system, where $J$ is the moment map of an effective Hamiltonian $S^{1}$-action on $M$. Let $\Sigma\subset M^{S^{1}}$ be a fixed surface. Then any rank one singular point $p\in\Sigma$ is of elliptic-regular type. Moreover, a rank zero singular point $p\in\Sigma$ is of elliptic-elliptic type (elliptic-hyperbolic type respectively) if and only if $p$ is an elliptic (hyperbolic resp.) singular point of $H|_{\Sigma}$ in the Morse sense.
4. Non-degeneracy of non-transition points
We will prove our main result in the remaining sections. Let $P\in\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$ be defined as in the statement of Theorem 1.1. In this section, we check the non-degeneracy of singular points in $\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})\setminus\{P\}$. We start by proving a helpful lemma.
@@ 4. Non-degeneracy of non-transition points
Fix positive numbers $c,r_{3},r_{4},r_{5}$ such that $\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold and $r_{3}>r_{5}$. For $t\in(0,1)$, define $H_{t}:\mathcal{P}(c,r_{3},r_{4},r_{5})\to{\mathbb{R}}$ by $H_{t}([(\rho_{0},\rho_{3},\rho_{4},\rho_{5})])=(1-t)|\rho_{4}+\rho_{5}|^{2}+t|\rho_{3}+\rho_{4}|^{2}$. Then $Q\in\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a singular point of $H_{t}$ if and only if there exists a representative $\rho$ with $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ and a nonzero number $a\in{\mathbb{R}}$ such that
(*) $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\textrm{ and }ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
@@
Consider a representative $\rho=(\rho_{0},\rho_{3},\rho_{4},\rho_{5})\in\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ of $Q$ such that $\rho_{0}=(c,0,0)$. By Lemma 2.4, $Q$ is a singular point if and only if there exists an $a\in{\mathbb{R}}$ such that $(a,0,0)\times\rho_{3}=t\rho_{4}\times\rho_{3}$ and $(a,0,0)\times\rho_{5}=(1-t)\rho_{4}\times\rho_{5}$. This implies that $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are coplanar, so we can further assume that $\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$. If $a=0$, then we have that $\rho_{3},\rho_{4},\rho_{5}$ are colinear. Since $\rho_{0}+\rho_{3}+\rho_{4}+\rho_{5}=0$, this implies that $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are all colinear, which is impossible since $\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold. Hence, we conclude that $Q$ is a singular point if and only if there exists a representative $\rho$ of the form $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ such that there exists an $a\in{\mathbb{R}}\setminus\{0\}$ with
$ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\textrm{ and }ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
(1) Regardless of the value of $t$, if $b_{3}=0$, then the equation $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$ implies that $b_{4}=0$. However, since $\rho_{0}+\rho_{3}+\rho_{4}+\rho_{5}=0$, this further implies that $b_{5}=0$ and thus $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are all colinear, which is impossible, so $b_{3}\neq 0$. Similarly, $b_{5}\neq 0$. Assume that $t\neq\dfrac{1}{2}$. If $b_{4}=0$, then Equation (*4.1) implies that $a=ta_{4},a=(1-t)a_{4}$, which is impossible unless $t=\dfrac{1}{2}$. Hence, we conclude that $b_{4}$ is nonzero whenever $t\neq\dfrac{1}{2}$.
(2) When $t\neq\dfrac{1}{2}$, we prove $K\neq-1$ by contradiction. If K=1K=-1, then $b_{5}=-\dfrac{1-t}{t}b_{3}$, so $b_{4}=-b_{3}-b_{5}=\dfrac{1-2t}{t}b_{3}=-\dfrac{1-2t}{1-t}b_{5}$. Since $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$ and $b_{3}\neq 0$, we can substitute $b_{4}$ by $\dfrac{1-2t}{t}b_{3}$ and divide $b_{3}$ on both sides to get that
@@
When $t=\dfrac{1}{2}$, $K=\dfrac{b_{5}}{b_{3}}$. Since $b_{3}+b_{4}+b_{5}=0$, we have $1+\dfrac{b_{4}}{b_{3}}+K=0$. Hence, if $b_{4}\neq 0$, $K\neq-1$. If $b_{4}=0$, then $b_{3}=-b_{5}$, so $a_{3}b_{5}-a_{5}b_{3}=b_{5}(a_{3}+a_{5})$. Since $r_{3}>r_{5}$, we have $|a_{3}|=\sqrt{r_{3}^{2}-b_{3}^{2}}>\sqrt{r_{5}^{2}-b_{5}^{2}}=|a_{5}|$ and thus $a_{3}b_{5}-a_{5}b_{3}=b_{5}(a_{3}+a_{5})\neq 0$. ∎
The following proposition takes care of almost all singular points except for the transition point and singular points that have the same moment image as the transition point.
@@
Let $r_{1},\ldots,r_{5}$ be positive numbers such that
  • $r_{4}<r_{5}<r_{3}$,
  • $|r_{1}+r_{2}-r_{3}|<r_{5}-r_{4}$,
  • $|r_{3}-r_{4}-r_{5}|<|r_{1}-r_{2}|<r_{3}+r_{4}-r_{5}$.
Let $M:=\mathcal{P}(r_{1},r_{2},r_{3},r_{4},r_{5})$. Consider the family $(J,H_{t})=(\ell_{12},t\ell_{34}^{2}+(1-t)\ell_{45}^{2})$ on $M$, where $t\in(0,1)$. Fix $|r_{1}-r_{2}|\leq c\leq r_{1}+r_{2}$ such that $c\neq r_{3}+r_{4}-r_{5}$. Let $M^{\textrm{red},c}:=J^{-1}(c)/S^{1}$ be the symplectic reduction at level $c$ and let $H^{\textrm{red},c}_{t}$ be the function to which $H_{t}$ descends. Then $H^{\textrm{red},c}_{t}$ is a Morse function with only elliptic singularities.
@@
By Lemma 2.5, $M^{\textrm{red},c}$ is a smooth manifold if $c\neq|r_{1}-r_{2}|,r_{1}+r_{2},r_{3}+r_{4}-r_{5}$. Moreover, if $c=|r_{1}-r_{2}|$ or $c=r_{1}+r_{2}$, then $M^{\textrm{red},c}=J^{-1}(c)/S^{1}=J^{-1}(c)$ is a smooth surface fixed by the $S^{1}$-action. Hence, by Proposition 2.2, $M^{\textrm{red},c}\cong\mathcal{P}(c,r_{3},r_{4},r_{5})$.
Let $Q\in M^{\textrm{red},c}$ be a singular point of $H^{\textrm{red},c}_{t}$. By Lemma 4.1, there exists a representative $\rho$ of $Q$ with $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ and a nonzero number $a\in{\mathbb{R}}$ such that
(*) $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\textrm{ and }ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
To compute the Hessian at $Q$, we find $W_{1},W_{2}\in T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ such that $W_{1}],[W_{2}]$ form a basis of $T_{Q}M^{\textrm{red},c}$ and then find the matrix representation of the Hessian of $H_{t}^{\textrm{red},c}$ under this basis. Since $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ is a submanifold of $S^{2}_{c}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$, it suffices to find a coordinate chart of $S^{2}_{c}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$ around $\rho$. By item (1) of Lemma 4.1, we have $b_{3},b_{5}\neq 0$. Depending on whether $b_{4}$ is zero, we choose different charts.
@@
Let $V_{\rho}$ be the real span of $V_{1},V_{2},V_{3}$. Define $W_{1}=\frac{\partial}{\partial z_{3}}-\frac{\partial}{\partial z_{5}}$ and $W_{2}=(a_{5}b_{3}-a_{3}b_{5})\frac{\partial}{\partial y_{0}}-b_{3}b_{5}\frac{\partial}{\partial x_{3}}+b_{3}b_{5}\frac{\partial}{\partial x_{5}}$. It’s straightforward to verify that $W_{1},W_{2},V_{1},V_{2},V_{3}$ form a basis of $T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$. Since $T_{Q}M^{\textrm{red},c}\cong T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})/V_{\rho}$, $W_{1}],[W_{2}]$ form a basis of $T_{Q}M^{\textrm{red},c}$.
@@
Since $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ is a principal $SO(3)$-bundle over $\mathcal{P}(c,r_{3},r_{4},r_{5})$, $\operatorname{Hess}(H_{t}^{\textrm{red},c})_{Q}([W_{i}],[W_{j}])=\operatorname{Hess}(H_{t}^{\textrm{red},c})_{\rho}(W_{i},W_{j})$, and we will compute the latter by extending $W_{1},W_{2}$ locally on $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$. Let $\widetilde{W}_{i}$ be a local extension on $W_{i}$ for $i=1,2$. We can write it as
$\widetilde{W}_{i}=f^{i}\frac{\partial}{\partial y_{0}}+g^{i}\frac{\partial}{\partial z_{0}}+\sum_{j=3}^{5}f_{j}^{i}\frac{\partial}{\partial x_{j}}+\sum_{j=3}^{5}g_{j}^{i}\frac{\partial}{\partial z_{j}}$
Since $\widetilde{W}_{i}$ is tangent to $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$, the coefficient functions satisfy the following equations:
(1) $f^{i}\frac{y_{0}}{x_{0}}-g^{i}\frac{z_{0}}{x_{0}}+f_{3}^{i}+f_{4}^{i}+f_{5}^{i}=0$
(2) $f^{i}-\sum_{j=3}^{5}\frac{x_{j}}{y_{j}}f_{j}^{i}-\sum_{j=3}^{5}\frac{z_{j}}{y_{j}}g_{j}^{i}=0$
(3) $g^{i}+g_{3}^{i}+g_{4}^{i}+g_{5}^{i}=0$
@@
$W_{1}(\widetilde{W}_{i}H_{1}^{\textrm{red},c})(\rho)=\frac{a_{4}b_{3}-a_{3}b_{4}}{b_{3}}\left(\frac{\partial f_{3}^{i}}{\partial z_{3}}(\rho)-\frac{b_{3}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{3}}(\rho)-\frac{\partial f_{3}^{i}}{\partial z_{5}}(\rho)+\frac{b_{3}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{5}}(\rho)\right)-g_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}$
$W_{1}(\widetilde{W}_{i}H_{0}^{\textrm{red},c})(\rho)=\frac{a_{4}b_{5}-a_{5}b_{4}}{b_{5}}\left(\frac{\partial f_{5}^{i}}{\partial z_{3}}(\rho)-\frac{b_{5}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{3}}(\rho)-\frac{\partial f_{5}^{i}}{\partial z_{5}}(\rho)+\frac{b_{5}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{5}}(\rho)\right)+g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}}$
Since $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$, $ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4})$ and $b_{3}+b_{4}+b_{5}=0$, we have
$W_{1}(\widetilde{W}_{i}H_{t}^{\textrm{red},c})(\rho)$ $=tW_{1}(\widetilde{W}_{i}H_{1}^{\textrm{red},c})(\rho)+(1-t)W_{1}(\widetilde{W}_{i}H_{0}^{\textrm{red},c})(\rho)$
$=a\left(\frac{\partial(f_{3}^{i}+f_{4}^{i}+f_{5}^{i})}{\partial z_{3}}(\rho)-\frac{\partial(f_{3}^{i}+f_{4}^{i}+f_{5}^{i})}{\partial z_{5}}(\rho)\right)-tg_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}+(1-t)g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}}$

### v2

@@ 3.2.2. Rank one points
Recall that a point $p\in M$ has rank one if $dJ_{p}$ and $dH_{p}$ span a one-dimensional vector space. When $dJ_{p}\neq 0$, $p$ is not a fixed point of the $S^{1}$-action. If $S^{1}$ acts freely on $p$, we can apply symplectic reduction (at least locally) at the level $J(p)$ and study the type of singularity of $p]\in M^{\textrm{red},J(p)}=J^{-1}(J(p))/S^{1}$ of the reduced function $H^{\textrm{red},J(p)}$ to which $H$ descends.
@@
[LP24, Lemma 2.6] Let $(M,\omega,F=(J,H))$ be a 4-dimensional integrable system, where $J$ is the moment map of an effective Hamiltonian $S^{1}$-action on $M$. Let $p\in M$ be a point on which $S^{1}$ acts freely. Let $j=J(p)$. Then $p$ is a rank one singular point of $F$ if and only if $[p]$ is a singular point of $H^{\textrm{red},j}$. Moreover, $p$ is a non-degenerate singular point of $F$ if and only if $[p]$ is non-degenerate for $H^{\textrm{red},j}$ (in the Morse sense) and the type of $p$ is elliptic-regular (respectively hyperbolic-regular) if and only if $[p]$ is an elliptic (respectively hyperbolic) singular point of $H^{\textrm{red},j}$.
@@
The proof of the first claim is similar to that of Lemma 3.5, so we leave it to the readers to verify. Since $X_{J}(0)=0$, the real span of $X_{J}(0)$ and $X_{H}(0)$ is the same as the real span of $X_{H}(0)$. Let $L$ denote this real span. Then $L^{\omega}={\mathbb{R}}(\frac{\partial}{\partial z_{1}},\frac{\partial}{\partial\overline{z}_{1}},X_{H}(0))$. Hence, we can use the equivalent classes of $\frac{\partial}{\partial z_{1}}$ and $\frac{\partial}{\partial\overline{z}_{1}}$ as a basis of $L^{\omega}/L$. Under this basis, we have the following matrix representation:
$\Omega=\begin{pmatrix}0&\frac{i}{2}\\ -\frac{i}{2}&0\end{pmatrix}\quad\text{and}\quad\operatorname{Hess}(J)=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$.
Hence, $A_{1,0}=\Omega^{-1}\operatorname{Hess}(J)$ has eigenvalues $\pm 2i$, so by definition $(0,0)$ is an elliptic-regular point of $(J,H)$. ∎
@@
Let $(M,\omega,(J,H))$ be a 4-dimensional integrable system, where $J$ is the moment map of an effective Hamiltonian $S^{1}$-action on $M$. Let $\Sigma\subset M^{S^{1}}$ be a fixed surface. Then any rank one singular point $p\in\Sigma$ is of elliptic-regular type. Moreover, a rank zero singular point $p\in\Sigma$ is of elliptic-elliptic type (elliptic-hyperbolic type, respectively) if and only if $p$ is an elliptic (hyperbolic resp.) singular point of $H|_{\Sigma}$ in the Morse sense.
4. Morse Functions on quadrilateral spaces
Recall that Lemma 3.7 implies that the non-degeneracy of rank one singularities can be checked by passing to the symplectic reduced spaces. We implement this idea in this section by studying functions on quadrilateral spaces, which, by Proposition 2.2, are symplectomorphic to the reduced spaces of interest. The main result, Proposition 4.2, combined with the aforementioned lemma and proposition, will establish the non-degeneracy of singularities of $F_{t}$. We start by proving a key lemma.
@@ 4. Morse Functions on quadrilateral spaces
Let $c,r_{3},r_{4},r_{5}$ be positive numbers such that $\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold. For $t\in(0,1)$, define $H_{t}:\mathcal{P}(c,r_{3},r_{4},r_{5})\to{\mathbb{R}}$ by $H_{t}([(\rho_{0},\rho_{3},\rho_{4},\rho_{5})])=(1-t)|\rho_{4}+\rho_{5}|^{2}+t|\rho_{3}+\rho_{4}|^{2}$. Then $m\in\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a singular point of $H_{t}$ if and only if there exists a representative $\rho$ with $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ and a nonzero number $a\in{\mathbb{R}}$ such that
(*) $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\quad\textrm{ and }\quad ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
@@
Consider a representative $\rho=(\rho_{0},\rho_{3},\rho_{4},\rho_{5})\in\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ of $m$ such that $\rho_{0}=(c,0,0)$. By Lemma 2.4, $m$ is a singular point of $H_{t}$ if and only if there exists $a\in{\mathbb{R}}$ such that $(a,0,0)\times\rho_{3}=t\rho_{4}\times\rho_{3}$ and $(a,0,0)\times\rho_{5}=(1-t)\rho_{4}\times\rho_{5}$. This implies that $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are coplanar, so we can further assume that $\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$. If $a=0$, then $\rho_{3},\rho_{4},\rho_{5}$ are colinear. Since $\rho_{0}+\rho_{3}+\rho_{4}+\rho_{5}=0$, this implies that $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are all colinear, which is impossible since $\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold. Hence, we conclude that $m$ is a singular point of $H_{t}$ if and only if there exists a representative $\rho$ of the form $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ such that there exists $a\in{\mathbb{R}}\setminus\{0\}$ with
$ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\quad\textrm{ and }\quad ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
(1) If $b_{3}=0$, then the equation $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$ implies that $b_{4}=0$. However, since $\rho_{0}+\rho_{3}+\rho_{4}+\rho_{5}=0$, this further implies that $b_{5}=0$ and thus $\rho_{0},\rho_{3},\rho_{4},\rho_{5}$ are all colinear, which is impossible, so $b_{3}\neq 0$. Similarly, $b_{5}\neq 0$. If $b_{4}=0$, then Equation (*4.1) implies that $a=ta_{4}$ and $a=(1-t)a_{4}$. Since both $a$ and $a_{4}$ are nonzero, $t=\frac{1}{2}$. Hence, we conclude that $b_{4}$ is nonzero whenever $t\neq\dfrac{1}{2}$.
(2) When $t\neq\dfrac{1}{2}$, we prove $K\neq-1$ by contradiction. If K=1K=-1, then $b_{5}=-\dfrac{1-t}{t}b_{3}$, so $b_{4}=-b_{3}-b_{5}=\dfrac{1-2t}{t}b_{3}=-\dfrac{1-2t}{1-t}b_{5}$. Since $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$ and $b_{3}\neq 0$, we can substitute $b_{4}$ by $\dfrac{1-2t}{t}b_{3}$ and divide by $b_{3}$ on both sides to get
@@
When $t=\dfrac{1}{2}$, $K=\dfrac{b_{5}}{b_{3}}$. Since $b_{3}+b_{4}+b_{5}=0$, we have $1+\dfrac{b_{4}}{b_{3}}+K=0$. Hence, if $b_{4}\neq 0$, $K\neq-1$. If $b_{4}=0$, then $b_{3}=-b_{5}$, so $a_{3}b_{5}-a_{5}b_{3}=b_{5}(a_{3}+a_{5})$. Assume for the sake of contradiction that $b_{5}(a_{3}+a_{5})=0$, then $a_{3}=-a_{5}$, because $b_{5}$ is nonzero. This implies $r_{3}=r_{5}$. Furthermore, since $c+a_{3}+a_{4}+a_{5}=0$, we have $c+a_{4}=0$. Since $b_{4}=0$, $|a_{4}|=r_{4}$, so $c=r_{4}$. However, this means that $c+r_{3}-r_{4}-r_{5}=0$, contradicting the assumption that $\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold. Hence, we conclude that $a_{3}b_{5}-a_{5}b_{3}\neq 0$. ∎
@@
Let $c,r_{3},r_{4},r_{5}$ be positive numbers such that $M:=\mathcal{P}(c,r_{3},r_{4},r_{5})$ is a smooth manifold. For $t\in(0,1)$, define $H_{t}:\mathcal{P}(c,r_{3},r_{4},r_{5})\to{\mathbb{R}}$ by $H_{t}([(\rho_{0},\rho_{3},\rho_{4},\rho_{5})])=(1-t)|\rho_{4}+\rho_{5}|^{2}+t|\rho_{3}+\rho_{4}|^{2}$. Then $H_{t}$ is a Morse function with only elliptic singularities.
@@
Let $m\in M$ be a singular point of $H_{t}$. By Lemma 4.1, there exists a representative $\rho$ of $m$ with $\rho_{0}=(c,0,0),\rho_{i}=(a_{i},b_{i},0)$ for $i=3,4,5$ and a nonzero number $a\in{\mathbb{R}}$ such that
(*) $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})\quad\textrm{ and }\quad ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4}).$
To compute the Hessian at $m$, we find $W_{1},W_{2}\in T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ such that $W_{1}],[W_{2}]$ form a basis of $T_{m}M$ and then find the matrix representation of the Hessian of $H_{t}$ under this basis. Since $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ is a submanifold of $S^{2}_{c}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$, it suffices to find a coordinate chart of $S^{2}_{c}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}$ around $\rho$. By item (1) of Lemma 4.1, we have $b_{3},b_{5}\neq 0$. Depending on whether $b_{4}$ is zero, we choose different charts.
@@
Let $V_{\rho}$ be the real span of $V_{1},V_{2},V_{3}$. Define $W_{1}=\frac{\partial}{\partial z_{3}}-\frac{\partial}{\partial z_{5}}$ and $W_{2}=(a_{5}b_{3}-a_{3}b_{5})\frac{\partial}{\partial y_{0}}-b_{3}b_{5}\frac{\partial}{\partial x_{3}}+b_{3}b_{5}\frac{\partial}{\partial x_{5}}$. It is straightforward to verify that $W_{1},W_{2},V_{1},V_{2},V_{3}$ form a basis of $T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ which is the kernel of $d\mu_{\rho}$, where $\mu:S^{2}_{c}\times S^{2}_{r_{3}}\times S^{2}_{r_{4}}\times S^{2}_{r_{5}}\to{\mathbb{R}}^{3}$ is the moment map of the $SO(3)$-action. Since $T_{m}M\cong T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})/V_{\rho}$, $W_{1}],[W_{2}]$ form a basis of $T_{m}M$.
@@
Since $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$ is a principal $SO(3)$-bundle over $\mathcal{P}(c,r_{3},r_{4},r_{5})$, $\operatorname{Hess}(H_{t})_{m}([W_{i}],[W_{j}])=\operatorname{Hess}(H_{t})_{\rho}(W_{i},W_{j})$. We will compute the latter by extending $W_{1},W_{2}$ locally on $\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})$. Let $\widetilde{W}_{i}$ be a local extension of $W_{i}$ for $i=1,2$. We can write it as
$\widetilde{W}_{i}=f^{i}\frac{\partial}{\partial y_{0}}+g^{i}\frac{\partial}{\partial z_{0}}+\sum_{j=3}^{5}f_{j}^{i}\frac{\partial}{\partial x_{j}}+\sum_{j=3}^{5}g_{j}^{i}\frac{\partial}{\partial z_{j}}.$
Since $\widetilde{W}_{i}\in T_{\rho}\widetilde{\mathcal{P}}(c,r_{3},r_{4},r_{5})=\textrm{ker}(d\mu_{\rho})$, the coefficient functions satisfy the following equation:
(1) $f^{i}\frac{y_{0}}{x_{0}}-g^{i}\frac{z_{0}}{x_{0}}+f_{3}^{i}+f_{4}^{i}+f_{5}^{i}=0.$
@@
$W_{1}(\widetilde{W}_{i}H_{1})(\rho)=\frac{a_{4}b_{3}-a_{3}b_{4}}{b_{3}}\left(\frac{\partial f_{3}^{i}}{\partial z_{3}}(\rho)-\frac{b_{3}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{3}}(\rho)-\frac{\partial f_{3}^{i}}{\partial z_{5}}(\rho)+\frac{b_{3}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{5}}(\rho)\right)-g_{3}^{i}(\rho)\frac{b_{4}}{b_{3}},$
$W_{1}(\widetilde{W}_{i}H_{0})(\rho)=\frac{a_{4}b_{5}-a_{5}b_{4}}{b_{5}}\left(\frac{\partial f_{5}^{i}}{\partial z_{3}}(\rho)-\frac{b_{5}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{3}}(\rho)-\frac{\partial f_{5}^{i}}{\partial z_{5}}(\rho)+\frac{b_{5}}{b_{4}}\frac{\partial f_{4}^{i}}{\partial z_{5}}(\rho)\right)+g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}}.$
Since $ab_{3}=t(a_{4}b_{3}-a_{3}b_{4})$, $ab_{5}=(1-t)(a_{4}b_{5}-a_{5}b_{4})$, and $b_{3}+b_{4}+b_{5}=0$, we have
$W_{1}(\widetilde{W}_{i}H_{t})(\rho)$ $=tW_{1}(\widetilde{W}_{i}H_{1})(\rho)+(1-t)W_{1}(\widetilde{W}_{i}H_{0})(\rho)$
$=a\left(\frac{\partial(f_{3}^{i}+f_{4}^{i}+f_{5}^{i})}{\partial z_{3}}(\rho)-\frac{\partial(f_{3}^{i}+f_{4}^{i}+f_{5}^{i})}{\partial z_{5}}(\rho)\right)-tg_{3}^{i}(\rho)\frac{b_{4}}{b_{3}}+(1-t)g_{5}^{i}(\rho)\frac{b_{4}}{b_{5}}.$